A 20% ACL solution and a 70% HCL solution are to be mixed to obtain 900ml of a 30% HCL solution. How many milliliters of solution must be mixed?
Answer by rfer(10432)

.2x+.7(900-x)=900*.3
.2x+630-.7x=270
-.5x=-360
x=720 ml of 20%
900-x=180 ml 70%